# Categories of Line Defects and Cohomological Hall Algebras

### v2

@@ A.2.1 Three-node-one-loop quiver
$\begin{array}[]{ccc}\varphi^{1\Leftarrow 1}(u)=1&\varphi^{1\Leftarrow 2}(u)=u+h_{1}&\varphi^{1\Leftarrow 3}(u)=\frac{1}{u-h_{3}}\\ \varphi^{2\Leftarrow 1}(u)=\frac{1}{u-h_{1}}&\varphi^{2\Leftarrow 2}(u)=1&\varphi^{2\Leftarrow 3}(u)=u+h_{2}\\ \varphi^{3\Leftarrow 1}(u)=u+h_{3}&\varphi^{3\Leftarrow 2}(u)=\frac{1}{u-h_{2}}&\varphi^{3\Leftarrow 3}(u)=1\,.\end{array}$ (469)
Now consider the framing |1||$\rightarrow$ 1|| ∗ → 1 |=1. Let’s compute the state and its eigenvalue of $\psi^{(i)}(z)$ order by order. By construction, these eigenvalues are
@@
    Level-$1$. After adding the this atom, the state has charge function
@@
    Level-$2$. After adding the this atom, the state has charge function
@@
    Level-$2$. Applying $e^{(i)}_{n}$ on $e^{(1)}_{0}|0\rangle_{1}$, after imposing the quadratic relations and (476), we have
@@
Now repeat this argument for the other two framings of this kind, i.e. |2|=1|$\rightarrow$ 2|=1| ∗ → 2 | = 1 or |3|=1|$\rightarrow$ 3|=1| ∗ → 3 | = 1, we obtain
@@ Appendix B BPS particles and $(2,2)$ Supersymmetric Quantum Mechanics
i.e. we identify $dx=\psi^{-}_{-}$ and $d\bar{z}=\psi^{-}_{+}$. The $\psi^{+}_{\beta}$ are canonically conjugate variables to these. The states $|\alpha\rangle$ are 1-forms, while $|A\rangle$ are 00- and $2$-forms.
@@ B.1.3 Combining the traps
An $\Omega$ deformation in the holomorphic plane is a well-understood way to trap the 4D dynamics near the origin of $\mathbb{C}$. The $\Omega$ deformation in a plane effectively preserves a $(2,2)$ 2D sub-algebra of the 4D SUSY algebra. The 2D $(2,2)$ super-algebra has central charges which receive contributions both from the 4D central charge and from the angular momentum $J$ in the holomorphic plane. The angular momentum contribution will appear as a shift of $H$: the Hamiltonian is not $Q$-exact but $H-J$ still is. Conversely, we can define the $\Omega$ deformation by presenting the 4D system as a 2D system with $(2,2)$ SUSY and turning on a “twisted mass” for the rotational symmetry in the holomorphic plane.
@@ B.1.4 Algebras of (trapped) BPS particles
In particular, if some no-exotic statement implies that $A$ is supported in ghost number 00, we can place an empty trap in the middle of $A\otimes A^{!}\otimes A$, we get a map $A\otimes A\to A$ which should be associative, making $A$ into an algebra. Then the remaining relations do resemble axioms satisfied by $A^{!}$. [^30] In examples of Koszul duality, $A\otimes A^{!}$ equipped with a canonical differential is quasi-isomorphic to $\mathbb{C}$ only up to a degree shift. $A\otimes(A^{!})^{\vee}$ would be a better candidate.
@@ B.2 Quiver quantum mechanics, equivariant cohomology and CoHA
The conventional route to the CoHA is to consider the SQM associated to a quiver Denef:2002ru . This is a dimensional reduction of a 4d non-Abelian SUSY gauge theory: the fermions which used to be chiral or anti-chiral before the dimensional reduction now couple differently to the electric field $P_{\alpha\beta}$ and the magnetic field $[X^{\alpha\gamma},X^{\delta\beta}]\epsilon_{\gamma\delta}$. Physically, this is due to the fact that the quiver arises from a D-brane description of the BPS particles and string theory/supergravity does not have the $SU(2)_{R}$ symmetry. The discussion above will have to be adjusted accordingly.

### v3

@@ A.2.1 Three-node-one-loop quiver
$\begin{array}[]{ccc}\varphi^{1\Leftarrow 1}(u)=1\qquad&\varphi^{1\Leftarrow 2}(u)=u+h_{1}\qquad&\varphi^{1\Leftarrow 3}(u)=\frac{1}{u-h_{3}}\\ \varphi^{2\Leftarrow 1}(u)=\frac{1}{u-h_{1}}\qquad&\varphi^{2\Leftarrow 2}(u)=1\qquad&\varphi^{2\Leftarrow 3}(u)=u+h_{2}\\ \varphi^{3\Leftarrow 1}(u)=u+h_{3}\qquad&\varphi^{3\Leftarrow 2}(u)=\frac{1}{u-h_{2}}\qquad&\varphi^{3\Leftarrow 3}(u)=1\,.\end{array}$ (469)
Now consider the framing $|*\rightarrow 1|$=1. Let’s compute the state and its eigenvalue of $\psi^{(i)}(z)$ order by order. By construction, these eigenvalues are
@@
    Level-11. After adding the this atom, the state has charge function
@@
    Level-22. After adding the this atom, the state has charge function
@@
    Level-22. Applying $e^{(i)}_{n}$ on $e^{(1)}_{0}|0\rangle_{1}$, after imposing the quadratic relations and (476), we have
@@
Now repeat this argument for the other two framings of this kind, i.e. $|*\rightarrow 2|=1$ or $|*\rightarrow 3|=1$, we obtain
@@ Appendix B BPS particles and $(2,2)$ Supersymmetric Quantum Mechanics
i.e. we identify $dx=\psi^{-}_{-}$ and $d\bar{z}=\psi^{-}_{+}$. The $\psi^{+}_{\beta}$ are canonically conjugate variables to these. The states $|\alpha\rangle$ are 1-forms, while $|A\rangle$ are 0- and $2$-forms.
@@ B.1.3 Combining the traps
An $\Omega$ deformation in the holomorphic plane is a well-understood way to trap the 4D dynamics near the origin of $\mathbb{C}$. The $\Omega$ deformation in a plane effectively preserves a $(2,2)$ 2D sub-algebra of the 4D SUSY algebra. The 2D $(2,2)$ super-algebra has central charges which receive contributions both from the 4D central charge and from the angular momentum $J$ in the holomorphic plane. The angular momentum contribution will appear as a shift of $H$: the Hamiltonian is not $Q$-exact but HJH-J still is. Conversely, we can define the $\Omega$ deformation by presenting the 4D system as a 2D system with $(2,2)$ SUSY and turning on a “twisted mass” for the rotational symmetry in the holomorphic plane.
@@ B.1.4 Algebras of (trapped) BPS particles
In particular, if some no-exotic statement implies that $A$ is supported in ghost number 0, we can place an empty trap in the middle of $A\otimes A^{!}\otimes A$, we get a map $A\otimes A\to A$ which should be associative, making $A$ into an algebra. Then the remaining relations do resemble axioms satisfied by $A^{!}$. [^32] In examples of Koszul duality, $A\otimes A^{!}$ equipped with a canonical differential is quasi-isomorphic to $\mathbb{C}$ only up to a degree shift. $A\otimes(A^{!})^{\vee}$ would be a better candidate.
@@ B.2 Quiver quantum mechanics, equivariant cohomology and CoHA
The conventional route to the CoHA is to consider the SQM associated to a quiver Denef:2002ru . This is a dimensional reduction of a 4d non-Abelian SUSY gauge theory: the fermions which used to be chiral or anti-chiral before the dimensional reduction now couple differently to the electric field $P_{\alpha\beta}$ and the magnetic field $X^{\alpha\gamma},X^{\delta\beta}]\epsilon_{\gamma\delta}$. Physically, this is due to the fact that the quiver arises from a D-brane description of the BPS particles and string theory/supergravity does not have the $SU(2)_{R}$ symmetry. The discussion above will have to be adjusted accordingly.